Investing 20000 part at 4% and part at 6%.If the annual income from investments was equal how much was invested at each rate
Answer by rfer(16322)

0.04x=0.06(20000-x)
0.04x=1200-0.06x
0.1x=1200
x=12000 at 4%
20000-x=8000 at 6%